Martin asked his grandfather how old he was, he replied “my age is 4 times your age, but 5 years ago it was 5 times yours.” How old is his grandfather?
Answer by ikleyn(52786)

.
Martin asked his grandfather how old he was, he replied “my age is 4 times your age, but 5 years ago it was 5 times yours.” How old is his grandfather?
~~~~~~~~~~~~~~~~~~~~~~~~~~~


Let x be the Martin's age.

Then the grandfather's current age is  4x, according to the problem.


5 years ago

    5(x-5) = 4x-5.     (5 years ago the grandfather's age was 5 times Martin's age)


Simplify and find x

    5x - 25 = 4x - 5

    5x - 4x = 25 - 5

       x    =   20.


ANSWER.  Grandfather is 4*20 = 80 years old.

Solved.

-------------------

You may check it on your own that my answer satisfies all problem descriptions.
(I did it mentally for myself. Checking is an integral part of any solution, as you know or should know).